Calculate the nuclear binding energy of 55/25 Mn in joules.

Constants and conversion factors
The atomic mass of 55/25 Mn is 54.938 .
The speed of light is c= 3.00×10^8 M/S .
The mass of a proton is 1.0078 amu .
The mass of a neutron is 1.0087 amu .
The mass of an electron is 5.4858×10^-4 amu .
. 1kg=6.022*10^26 amu
. 1 J= 1 Kg* m^2/s^2
. 1 MeV= 1.602*10^-13 Joules
Answer #1
It may help to think of binding energy as "unbinding energy" in the sense that it is the energy required to split a nucleus apart into its constituentnucleons.

To solve the question calculate how much all the nucleons will weigh when apart, i.e.
Mass of:
25 protons = 25(1.0073)
25 protons = 25.1825 amu

25 neutrons = (55-25)(1.0087)
25 neutrons = 30.261 amu

Total mass of nucleons = 30.261+25.1825
Total mass of nucleons = 55.4435 amu

Now take the difference of the mass of nucleons and the mass of the Mn nucleus
55.4435-54.938=0.5055 amu

This difference in mass is also known as the mass defect of a nucleus. Remember, the mass of the constituent nucleons will always be greater than theobserved mass of the nucleus. This is because mass is lost in the form of energy when the nucleons come together to form the nucleus.

Now to calculate the binding energy, use the very famous formula E=mc^2 ;)

The binding energy of the nucleon will equal to the mass defect times the speed of light squared.
However, the formula works in SI units, so the mass defect has to be converted to kg
0.5055 amu = (0.5055)/(6.022x10^26)
0.5055 amu = 8.3942x10^-28

Now apply the formula,
E=?mc^2
E=(8.3942x10^-28)(3x10^8)^2
E=7.55x10^-11 J <----Final
Answer #2
It may help to think of binding energy as "unbinding energy" in the sense that it is the energy required to split a nucleus apart into its constituentnucleons.

To solve the question calculate how much all the nucleons will weigh when apart, i.e.
Mass of:
25 protons = 25(1.0073)
25 protons = 25.1825 amu

25 neutrons = (55-25)(1.0087)
25 neutrons = 30.261 amu

Total mass of nucleons = 30.261+25.1825
Total mass of nucleons = 55.4435 amu

Now take the difference of the mass of nucleons and the mass of the Mn nucleus
55.4435-54.938=0.5055 amu

This difference in mass is also known as the mass defect of a nucleus. Remember, the mass of the constituent nucleons will always be greater than theobserved mass of the nucleus. This is because mass is lost in the form of energy when the nucleons come together to form the nucleus.

Now to calculate the binding energy, use the very famous formula E=mc^2 ;)

The binding energy of the nucleon will equal to the mass defect times the speed of light squared.
However, the formula works in SI units, so the mass defect has to be converted to kg
0.5055 amu = (0.5055)/(6.022x10^26)
0.5055 amu = 8.3942x10^-28

Now apply the formula,
E=?mc^2
E=(8.3942x10^-28)(3x10^8)^2
E=7.55x10^-11 J <----Final
answered by: Henry N
Answer #3
the answer's absolutely right .
Check in other units.. u might find what u'r looking for
answered by: Dr Ajit Handa